PROSPECTS FOR THE FORMATION OF ULTRACOLD NaRb MOLECULES IN THE GROUND X^{1}Σ^{+}(v = 0; J = 0) STATE BY PHOTOASSOCIATION THROUGH THE INTERMEDIATE A^{1}Σ^{+} ∼ b^{3}Π COMPLEX

\underline{A.V. STOLYAROV}, V.I. PUPYSHEV, E.A. PAZYUK, A. ZAITSEVSKII, DEPARTMENT OF CHEMISTRY, MOSCOW STATE UNIVERSITY, GSP-2 LENINSKIE GORY 1/3, MOSCOW 119992, RUSSIA; O. DOCENKO, M. TAMANIS, R. FERBER, DEPARTMENT OF PHYSICS AND INSTITUTE OF ATOMIC PHYSICS AND SPECTROSCOPY, UNIVERSITY OF LATVIA, 19 RAINIS BOULEVARD, RIGA LV-1586, LATVIA; A. PASHOV, DEPARTMENT OF PHYSICS, SOFIA UNIVERSITY, 5 JAMES BOURCHIER BLVD, 1164 SOFIA, BULGARIA; H. KNÖCKEL, E. TIEMANN, INSTITUT FÜR QUANTENOPTIK, GOTTFRIED WILHELM LEIBNIZ UNIVER-SITÄT HANNOVER, WELFENGARTEN 1, 30167 HANNOVER, GERMANY.

The direct deperturbation analysis of the singlet-triplet A^{1}Σ^{+} ∼ b^{3}Π complex of Na^{85}Rb and Na^{87}Rb isotopomers was performed in the framework of the inverted channel-coupling approach ^{a}. Besides of Born-Oppenheimer potential energy curves of the mutually perturbed states and the off-diagonal spin-orbit (SO) A ∼ b coupling function, the SO splitting of the b^{3}Π state was determined due to the pronounced electronic- rotation interaction between the b^{3}Π_{0} and b^{3}Π_{1} components observed for high J-levels. Overall 24 mass-invariant fitting parameters have been required to reproduce about 2300 experimental term values of both isotopomers with a standard deviation of 0.012 cm^{−1} which is consistent with the uncertainty of the Fourier transform spectra measurement. The experimental relative intensities in the D^{1}Π → A ∼ b spectra highlighted the breakdown of the conventional oscillation theorem. The strong intensity borrowing effect caused by the SO coupling and enhanced by the violation of the Franck-Condon principle was analyzed. The calculated A ∼ b−a, X transition probabilities were applied for simulation of the stimulated Raman processes a^{3}Σ^{+} → A ∼ b → X^{1}Σ^{+} which can lead to efficient formation of ultracold NaRb molecules in the ground level v_{X} = 0; J_{X} = 0. The quantitative transition moments predicted for favorable lambda schemes show that such experiments are feasible.

aO. Docenko, M. Tamanis, R. Ferber, E.A. Pazyuk, A. Zaitsevskii, A.V. Stolyarov, A. Pashov, H. Knöckel, E. Tiemann Phys.

Rev. A accepted, 2007, in print